# VoiceBridge: Designing Latent Bridge Models for General Speech Restoration at Scale

### v1

@@ Abstract
Bridge models have recently been explored for speech enhancement tasks such as denoising, dereverberation, and super-resolution, while these efforts are typically confined to a single task or small-scale datasets, with constrained general speech restoration (GSR) capability at scale. In this work, we introduce VoiceBridge, a GSR system rooted in latent bridge models (LBMs), capable of reconstructing high-fidelity speech at full-band (i.e., 48 kHz) from various distortions. By compressing speech waveform into continuous latent representations, VoiceBridge models the diverse LQ-to-HQ tasks (namely, low-quality to high-quality) in GSR with a single latent-to-latent generative process backed by a scalable transformer architecture. To better inherit the advantages of bridge models from the data domain to the latent space, we present an energy-preserving variational autoencoder, enhancing the alignment between the waveform and latent space over varying energy levels. Furthermore, to address the difficulty of HQ reconstruction from distinctively different LQ priors, we propose a joint neural prior, uniformly alleviating the reconstruction burden of LBM. At last, considering the key requirement of GSR systems, human perceptual quality, a perceptually aware fine-tuning stage is designed to mitigate the cascading mismatch in generation while improving perceptual alignment. Extensive validation across in-domain and out-of-domain tasks and datasets (e.g., refining recent zero-shot speech and podcast generation results) demonstrates the superior performance of VoiceBridge. Demo samples can be visited at: https://VoiceBridge-demo.github.io/.
@@ 1 Introduction
Tractable Schrödinger Bridge (SB) models have proven to faithfully reconstruct the target distribution from an informative prior, overcoming the limitation of the noise prior in diffusion models (Chen et al., 2021; 2023; Li et al., 2025b). Recently, their advantages have been extended to speech enhancement tasks (Jukić et al., 2024; Wang et al., 2024a; Han et al., 2025), such as denoising (Jukić et al., 2024), dereverberation, and super-resolution (Li et al., 2025a), where the low-quality (LQ) observation naturally provides indicative information for the high-quality (HQ) target. Although these efforts have proposed innovations spanning forward process design (Jukić et al., 2024), model parameterization (Richter et al., 2025), training objectives (Zhang et al., 2025a), and cascaded architectures (Han et al., 2025; Li et al., 2025b), their designs are usually confined to a single task or small-scale datasets, not yet achieving general speech restoration (GSR) at scale to address various real-world speech degradations.
In this work, we propose VoiceBridge, a general speech restoration (GSR) system (Benesty et al., 2006; Kolbæk et al., 2016; Das et al., 2021) rooted in a latent bridge model (LBM), where the diverse LQ-to-HQ tasks in GSR are modeled with a single latent-to-latent generative process backed by a transformer architecture. By encoding both the LQ and HQ speech signals into continuous representations in a small latent space, VoiceBridge aims to preserve the advantages of the data-to-data sampling nature of bridge models from data space to latent space while achieving efficient training. Towards generalizable modeling, VoiceBridge explores combining the advantages of transformer architecture, namely good scalability that has been verified in both image (Peebles & Xie, 2023; Bao et al., 2023) and audio (Evans et al., 2024a) generation, with the bridge generative framework, further strengthening GSR performance at scale.
To better inherit the advantages of bridge generative models on speech enhancement tasks in the data domain, e.g., bandwidth extension in the waveform domain (Li et al., 2025a), within LBMs, a natural idea is to preserve the characteristics in data space into latent representations. Motivated by this, we develop an energy-preserving variational autoencoder (EP-VAE) to achieve stronger consistency between the waveform and latent space than vanilla VAE (Kingma et al., 2019; Evans et al., 2024a). Specifically, we modify the VAE training objective, requiring a linear scaling operation in the latent space to be manifested in the decoded waveform space. In comparison with vanilla VAE that regularizes data reconstruction in a single scale, EP-VAE introduces reconstruction regularization over varying energy levels, thus improving the waveform-latent consistency and resulting in a more structural latent space (Kouzelis et al., 2025; Yao et al., 2025; Skorokhodov et al., 2025; Yu et al., 2024) for LBM modeling.
@@
Given the joint neural prior and the HQ target encoded in a structural latent space, we have been able to establish a bridge transformer to tackle the various restoration tasks at scale. Considering the key requirement of GSR systems (Manjunath, 2009; Manocha et al., 2022; Babaev et al., 2024), human perceptual quality, which is a main difference from task-specific SE systems pursuing accurate waveform reconstruction (Li et al., 2025a; Jukić et al., 2024), we introduce a perceptual-aware fine-tuning stage. This design complements the LBM which is typically trained to fit the target distribution from prior distribution without explicitly considering human perception quality. Concretely, we introduce two popular speech metrics, PESQ (Rix et al., 2001) and UTMOS (Saeki et al., 2022), into the post-training stage of VoiceBridge, which optimizes both the LBM sampling and the EP-VAE decoding, minimizing their cascading mismatches in inference and aligning the entire generation process with human perceptual quality.
@@
  • We design VoiceBridge, a LBM-based GSR system tackling diverse LQ-to-HQ tasks with a single latent-to-latent generative process backed a transformer architecture.
  • We propose EP-VAE to strengthen the consistency between the waveform and the latent spaces over varying energy levels, thereby inheriting the advantages of the bridge generative framework from the data domain into the latent space.
  • We develop joint neural prior and perceptual-aware fine-tuning, further reducing the reconstruction burden of LBM while aligning it with human perceptual quality.
  • Comprehensive validation across in-domain and out-of-domain tasks and datasets, including refining recent zero-shot speech and podcast generation results, shows the superiority of VoiceBridge in reconstructing high-fidelity 48 kHz speech from various distortions.
@@ 2 Related Work
General speech restoration. Previous GSR systems have exploited mapping-based methods (Liu et al., 2022), adversarial training (Babaev et al., 2024), masked generative models (Li et al., 2024; Zhang et al., 2025b) or pretraining (Wang et al., 2025), conditional diffusion models (Serrà et al., 2022; Scheibler et al., 2024; Dhyani et al., 2025), and flow-matching methods (Niu, 2024). In VoiceBridge, we design an LBM that extends the advantages of the data-to-data generative framework from specific SE tasks to GSR at scale and introduce three novel techniques that holistically improve generation results. We provide a more detailed introduction to related work in Appendix A.
@@ 3 VoiceBridge
In this section, we first present our designed LBM-based GSR system, and then introduce the three innovations leading to holistic improvement, namely EP-VAEjoint neural prior, and perceptual-aware generation, respectively.
@@ 3.1 Latent Bridge Transformer for GSR
where $\mathcal{T}$ is sampled from a range of speech degradation methods, including additive noise, reverberation, bandwidth limitation, clipping, vocal effects, and their random mixtures. Namely, for each ground-truth $\mathbf{x}_{0}$, VoiceBridge considers numerous different LQ prior $\mathbf{x}_{1}$, rather than considering a fixed degradation method as a single speech enhancement task (Jukić et al., 2024; Li et al., 2025a; b). A more detailed data construction process at the training stage is introduced in Appendix B.
Given constructed LQ and HQ waveform pairs $(\mathbf{x}_{0},\mathbf{x}_{1})$, we directly compress the speech waveform into latent representations, obtaining the latent target and prior, $\mathbf{z}_{0}\in\mathbb{R}^{c\times l}$ and $\mathbf{z}_{1}\in\mathbb{R}^{c\times l}$. In comparison with alternative speech compression representations, the waveform latent naturally preserves the data-to-data generative process for the LQ-to-HQ speech tasks without suffering from area removal (Kong et al., 2025), which is important to a range of tasks in GSR. For example, for the down-sampling degradation and its mixture with other degradation methods, GSR systems are usually required to generate the 48 kHz waveform from its 2 kHz or even 1 kHz version (Andreev et al., 2023; Ku et al., 2025). In these settings, the prior information contained in the spectrograms (Mandel et al., 2023; Kong et al., 2025) and the latent space of the mel-spectrogram (Liu et al., 2022) will be very limited because of their large-scale area removal, which may hinder the final restoration performance as mentioned by (Lim et al., 2018; Kong et al., 2025).
@@
where $\hat{\mathbf{z}}_{0,\varphi}$ denotes the latent predicted by the neural network with trainable parameters $\varphi$; $\mathbf{z}_{0},\mathbf{z}_{1}$ are the paired latent representations encoded from the HQ signal and its degraded LQ versions, respectively; $\mathbf{z}_{t}$ is constructed with Equation 2. For generalizable training, we explore combining the advantage of transformer architecture, namely good scalability as shown in diffusion-based both image (Peebles & Xie, 2023; Bao et al., 2023) and audio generation (Evans et al., 2024b), with the bridge generative framework. More discussion of tractable Schrödinger bridge models is provided in Appendix C. The details of network architecture are introduced in Appendix D.
@@ 3.2 Energy-Preserving VAE
As shown in Equation 3, the prior and target in VoiceBridge, $(\mathbf{z}_{0},\mathbf{z}_{1})$, are compressed from the corresponding waveform $(\mathbf{x}_{0},\mathbf{x}_{1})$. The compression network, namely VAE (Kingma et al., 2019), is employed to reduce the redundant information in waveform space to allow efficient training, while ensuring the reconstruction quality. Generally, the training of audio waveform VAE (Evans et al., 2024b) with network parameters $\theta$ optimizes the objectives on data-space reconstruction $\mathcal{L}_{\text{data}}$ and latent-space regularization $\mathcal{L}_{\text{latent}}$:
$\mathcal{L}_{\text{vae}}$ $=\mathcal{L}_{\text{data}}(\mathcal{D}_{\theta}(\mathcal{E}_{\theta}(\mathbf{x})),\mathbf{x})+\mathcal{L}_{\text{latent}}(\mathcal{E}_{\theta}(\mathbf{x}),\mathbf{z}_{\text{ref}}),$ (4)
where the term $\mathcal{L}_{\text{data}}$ is typically composed of reconstruction loss, adversarial loss, and feature matching loss between input signal $\mathbf{x}$ and reconstructed version $\mathcal{D}(\mathcal{E}(\mathbf{x}))$, while the term $\mathcal{L}_{\text{latent}}$ represents the KL divergence regularization of the latent $\mathbf{z}=\mathcal{E}(\mathbf{x})$ with a known prior distribution $\mathbf{z}_{\text{ref}}$e.g., standard Gaussian. In recent task-specific bridge-based speech enhancement works, the data-to-data generative framework effectively exploits the instructive information contained in the LQ observation in either waveform space (Li et al., 2025a) or STFT representations (Jukić et al., 2024; Kong et al., 2025; Han et al., 2025). Hence, when designing a latent-to-latent GSR system, we naturally explore preserving the advantages of bridge models from the data space to the encoded latent space. In VoiceBridge, to facilitate latent SB modeling, we introduce a novel technique, EP-VAE, to enhance the consistency between data and latent from the perspective of scale equivariance. Specifically, we add an EP constraint into the VAE training objective, expanding the data-space loss $\mathcal{L}_{\text{data}}$ in Equation 4 to continuous energy levels with $\mathcal{L}_{\text{data}}^{\text{ep}}$:
@@ 3.3 Prior Convergence
Figure 2: The tSNE visualization (Maaten & Hinton, 2008) of the prior latent before and after prior convergence from augmented VCTK (Veaux & Yamagishi, 2017). Note the difference between the scale of axes of the two figures.
@@
where $\mathcal{L}_{\text{latent}}$ has been the joint neural prior loss consisting of MSE loss and cosine similarity loss as introduced in Appendix D, and $\mathcal{L}_{\text{data}}^{\text{ep}}$ is the data-space loss in EP-VAE training modified for prior convergence. Namely, as shown in Equation 6, the encoder $\mathcal{E}^{\text{np}}$ is fine-tuned with changed objectives to converge different LQ priors to HQ target, rather than still learning the latent presentation for reconstruction as Equation 5. Moreover, the EP constraint is preserved. When reducing the distance between LQ priors $\mathbf{z}_{1}$ and the HQ target $\mathbf{z}_{0}$ to achieve $\mathbf{z}_{1}^{\text{np}}$, scale equivariance is simultaneously required, which inherits the EP property to measure the consistency at varying energy levels. As shown in Figure 2, this technique significantly closes the distance between different LQ priors and the HQ target in the latent space, and naturally reduces the burden of LBM modeling. A further quantitative analysis on the prior distance is shown in Appendix G
Table 1: Evaluation results on three GSR testsets, including a standard benchmark, a simulated testset whose training split is outside our training dataset for cross evaluation, and an in-the-wild dataset. Best is bolded, and second best is underlined. * denotes results taken from the original paper. VoiceBridge outperforms most baselines in nearly all metrics, demonstrating strong restoration ability in complex environments.
Voicefixer-GSR (Liu et al., 2022) Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Input 1.943 2.962 2.858 2.395 2.633 1.965 2.674 VoiceFixer 2.043 3.302 3.971 3.005 3.439 3.800 4.159 Resemble-Enhance 2.000 3.412 4.045 3.137 3.608 4.224 4.507 UniverSE++ 2.346 3.275 3.961 2.976 3.376 3.648 4.009 AnyEnhance - 3.406* 4.073* 3.136* - - 4.308* VoiceBridge 2.436 3.469 4.083 3.210 4.330 4.246 4.098 DNS-with-Reverb (Reddy et al., 2021a) Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Input 1.176 1.767 1.481 1.377 1.323 0.019-0.019 1.689 VoiceFixer 1.407 3.355 3.982 3.052 2.526 2.961 3.704 Resemble-Enhance 1.154 3.502 4.021 3.205 2.756 3.248 4.583 UniverSE++ 1.098 2.548 3.649 2.244 1.669 2.463 3.006 AnyEnhance - 3.500* 4.040* 3.204* - - 3.722* VoiceBridge 1.521 3.552 4.045 3.270 3.373 4.188 4.563 DNS-Real (Reddy et al., 2021a) Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Input - 2.985 2.510 2.212 1.940 1.483 2.160 VoiceFixer - 3.174 3.919 2.875 2.351 2.938 3.529 Resemble-Enhance - 3.395 3.993 3.100 2.767 3.298 4.329 UniverSE++ - 2.999 3.660 2.641 2.306 2.603 3.317 AnyEnhance - 3.488* 3.977* 3.161* - - - VoiceBridge - 3.447 3.985 3.166 4.197 4.034 4.497
3.4 Perceptual-Aware Generation

### v3

# VoiceBridge: General Speech Restoration with One-step Latent Bridge Models

@@ Abstract
Bridge models have been investigated in speech enhancement but are mostly single-task, with constrained general speech restoration (GSR) capability. In this work, we propose VoiceBridge, a one-step latent bridge model (LBM) for GSR, capable of efficiently reconstructing 48 kHz fullband speech from diverse distortions. To inherit the advantages of data-domain bridge models, we design an energy-preserving variational autoencoder, enhancing the waveform-latent space alignment over varying energy levels. By compressing waveform into continuous latent representations, VoiceBridge models various GSR tasks with a single latent-to-latent generative process backed by a scalable transformer. To alleviate the challenge of reconstructing the high-quality target from distinctively different low-quality priors, we propose a joint neural prior for GSR, uniformly reducing the burden of the LBM in diverse tasks. Building upon these designs, we further investigate bridge training objective by jointly tuning LBM, decoder and discriminator together, transforming the model from a denoiser to generator and enabling one-step GSR without distillation. Extensive validation across in-domain (e.g., denoising and super-resolution) and out-of-domain tasks (e.g., refining synthesized speech) and datasets demonstrates the superior performance of VoiceBridge. Demos: https://VoiceBridgedemo.github.io/.
Machine Learning, ICML, VoiceBridge, Latent Bridge Models, General Speech Restoration, Speech Enhancement, Speech Restoration, Speech Signal Processing, Diffusion Models, Generative Models, Schrodinger Bridge
@@ 1 Introduction
Tractable Schrödinger Bridge (SB) models have proven to faithfully reconstruct the target distribution from an informative prior, overcoming the limitation of the noise prior in diffusion models (Chen et al., 2021, 2023; Li et al., 2025b). Recently, their advantages have been extended to speech enhancement tasks (Jukić et al., 2024; Wang et al., 2024a; Han et al., 2025), such as denoising (Jukić et al., 2024), dereverberation, and super-resolution (Li et al., 2025a), where the low-quality (LQ) observation naturally provides indicative information for the high-quality (HQ) target. Although these efforts have proposed innovations spanning forward process design (Jukić et al., 2024), model parameterization (Richter et al., 2025), training objectives (Zhang et al., 2025a), and cascaded architectures (Han et al., 2025; Li et al., 2025b), their designs are usually confined to a single task, not yet achieving general speech restoration (GSR) at scale to address various real-world speech degradations.
In this work, we propose VoiceBridge, a one-step general speech restoration (GSR) system (Benesty et al., 2006; Kolbæk et al., 2016; Das et al., 2021) rooted in a latent bridge model (LBM), where the diverse LQ-to-HQ tasks in GSR are modeled with a single latent-to-latent generative process backed by a transformer architecture. By encoding both the LQ and HQ speech signals into continuous representations in a small latent space, VoiceBridge aims to preserve the advantages of the data-to-data sampling nature of bridge models from data space to latent space while achieving efficient training.
To better inherit the advantages of bridge generative models on speech enhancement tasks in the data domain within LBMs, a natural idea is to preserve the characteristics in data space into latent representations. Motivated by this, we develop an energy-preserving variational autoencoder (EP-VAE) to achieve stronger consistency between the waveform and latent space than vanilla VAE (Kingma et al., 2019). Specifically, we modify the VAE training objective, requiring a linear scaling operation in the latent space to be manifested in the decoded waveform space. In comparison with vanilla VAE that regularizes data reconstruction in a single scale, EP-VAE introduces reconstruction regularization over varying energy levels, naturally improving the waveform-latent consistency and resulting in a more structural latent space (Kouzelis et al., 2025; Yao et al., 2025; Skorokhodov et al., 2025; Yu et al., 2024) for LBM modeling.
@@
To mitigate cascading mismatches between the LBM and the decoder, we perform a post-training stage that jointly fine-tunes the LBM and the VAE decoder with the encoder fixed. Direct supervision in the data domain improves the realism of LBM predictions and calibrates the decoder to the LBM-induced latent distribution, while preserving the encoder’s learned latent geometry. Furthermore, We novelly incorporate adversarial and perceptual losses to align the one-step prediction with the full conditional distribution, converting the LBM from a multi-step denoiser into a one-step generator without distillation, achieving streaming-rate synthesis with state-of-the-art quality.
@@
  • We design VoiceBridge, a LBM-based GSR system tackling diverse LQ-to-HQ tasks with a single latent-to-latent generative process backed a transformer architecture, and integrating multiple innovations to the LBM system.
  • We propose EP-VAE to inherit the data-space advantages of bridge models by strengthening consistency between the waveform and latent across energy levels, and a joint neural prior to alleviate LQ-to-HQ generation burden by enhancing the prior distribution.
  • We propose a novel denoiser-to-generator post-training process, mitigating bridge-decoder mismatches, improving generation quality, and enabling one-step inference without distillation.
  • Comprehensive validation across in-domain and out-of-domain tasks and datasets, including refining recent zero-shot speech and podcast generation results, shows the superiority of VoiceBridge in efficiently reconstructing high-fidelity 48 kHz speech from various distortions.
@@ 2 Related Work
General speech restoration. Previous GSR systems have exploited mapping-based methods (Liu et al., 2022), adversarial training (Babaev et al., 2024), masked generative models (Li et al., 2024; Zhang et al., 2025b) or pretraining (Wang et al., 2025), conditional diffusion models (Serrà et al., 2022; Scheibler et al., 2024; Dhyani et al., 2025), and flow-matching methods (Niu, 2024; Liu et al., 2023a; Ku et al., 2025). In VoiceBridge, we design an LBM that extends the advantages of the data-to-data generative framework from specific SE tasks to GSR at scale and introduce three novel techniques that holistically improve generation results. We provide a more detailed introduction to related work in Appendix A.
@@ 3 VoiceBridge
In this section, we first present our designed LBM-based GSR system, and then introduce the three innovations leading to holistic improvement, namely EP-VAEjoint neural prior, and denoiser-to-generator alignment, respectively.
@@ 3.1 Latent Bridge Transformer for GSR
where $\mathcal{T}$ is sampled from a range of speech degradation methods, including additive noise, reverberation, bandwidth limitation, clipping, vocal effects, and their random mixtures. Namely, for each ground-truth $\mathbf{x}_{0}$, VoiceBridge considers numerous different LQ prior $\mathbf{x}_{1}$, rather than considering a fixed degradation method as a single speech enhancement task (Jukić et al., 2024; Li et al., 2025a, b). A more detailed data construction process at the training stage is introduced in Appendix B.
Given constructed LQ and HQ waveform pairs $(\mathbf{x}_{0},\mathbf{x}_{1})$, we directly compress the speech waveform into latent representations, obtaining the latent target and prior, $\mathbf{z}_{0}\in\mathbb{R}^{c\times l}$ and $\mathbf{z}_{1}\in\mathbb{R}^{c\times l}$. Prior works (Li et al., 2025a; Jukić et al., 2024) mainly conduct bridge modeling in the data space, either in the waveform or spectrogram domain, to handle single tasks such as speech super-resolution or speech enhancement. However, we find that models in these domains struggle in the complex GSR scenario and 48kHz settings. Directly performing SB in the data space takes great modeling and computational effort. Meanwhile, raw speech signals include massive redundant information at high frequencies, which can be effectively compressed by VAEs, resulting in a more compact latent space. Latent compression reduces sequence length by more than an order of magnitude, enabling a single 544M transformer LBM to handle all tasks. Details experiments proving the necessity of the latent SB design is shown in Appendix I
In comparison with alternative speech compression representations, the waveform latent naturally preserves the data-to-data generative process for the LQ-to-HQ speech tasks without suffering from area removal (Kong et al., 2025), which is important to a range of tasks in GSR. For example, for the down-sampling degradation and its mixture with other degradation methods, GSR systems are usually required to generate the 48 kHz waveform from its 2 kHz or even 1 kHz version (Andreev et al., 2023; Ku et al., 2025). In these settings, the prior information contained in the spectrograms (Mandel et al., 2023; Kong et al., 2025) and the latent space of the mel-spectrogram (Liu et al., 2022) will be very limited because of their large-scale area removal, which may hinder the final restoration performance as mentioned by (Lim et al., 2018; Kong et al., 2025).
@@
where $\hat{\mathbf{z}}_{0,\varphi}$ denotes the latent predicted by the neural network with trainable parameters $\varphi$; $\mathbf{z}_{0},\mathbf{z}_{1}$ are the paired latent representations encoded from the HQ signal and its degraded LQ versions, respectively; $\mathbf{z}_{t}$ is constructed with Equation 2. For generalizable training, we explore combining the advantage of transformer architecture, namely good scalability as shown in diffusion-based both image (Peebles and Xie, 2023; Bao et al., 2023) and audio generation (Evans et al., 2024b), with the bridge generative framework. More discussion of tractable Schrödinger bridge models is provided in Appendix C. The details of network architecture are introduced in Appendix D.
@@ 3.2 Energy-Preserving VAE
As shown in Equation 3, the prior and target in VoiceBridge, $(\mathbf{z}_{0},\mathbf{z}_{1})$, are compressed from the corresponding waveform $(\mathbf{x}_{0},\mathbf{x}_{1})$. The compression network, namely VAE (Kingma et al., 2019), is employed to reduce the redundant information in waveform space to allow efficient training, while ensuring the reconstruction quality. Generally, the training of audio waveform VAE (Evans et al., 2024b) with network parameters $\theta$ optimizes the objectives on data-space reconstruction $\mathcal{L}_{\text{data}}$ and latent-space regularization $\mathcal{L}_{\text{lat.}}$:
$\mathcal{L}_{\text{vae}}$ $=\mathcal{L}_{\text{data}}(\mathcal{D}_{\theta}(\mathcal{E}_{\theta}(\mathbf{x})),\mathbf{x})+\mathcal{L}_{\text{lat.}}(\mathcal{E}_{\theta}(\mathbf{x}),\mathbf{z}_{\text{ref}}),$ (4)
where the term $\mathcal{L}_{\text{data}}$ is typically composed of reconstruction loss, adversarial loss, and feature matching loss between input signal $\mathbf{x}$ and reconstructed version $\mathcal{D}(\mathcal{E}(\mathbf{x}))$, while the term $\mathcal{L}_{\text{lat.}}$ represents the KL divergence regularization of the latent $\mathbf{z}=\mathcal{E}(\mathbf{x})$ with a known prior distribution $\mathbf{z}_{\text{ref}}$e.g., standard Gaussian. In recent task-specific bridge-based speech enhancement works, the data-to-data generative framework effectively exploits the instructive information contained in the LQ observation in either waveform space (Li et al., 2025a) or STFT representations (Jukić et al., 2024; Kong et al., 2025; Han et al., 2025). Hence, when designing a latent-to-latent GSR system, we naturally explore preserving the advantages of bridge models from the data space to the encoded latent space. In VoiceBridge, to facilitate latent SB modeling, we introduce a novel technique, EP-VAE, to enhance the consistency between data and latent from the perspective of scale equivariance. Specifically, we add an EP constraint into the VAE training objective, expanding the data-space loss $\mathcal{L}_{\text{data}}$ in Equation 4 to continuous energy levels with $\mathcal{L}_{\text{data}}^{\text{ep}}$:
$\mathcal{L}_{\text{ep}}$ $=\mathcal{L}^{\text{ep}}_{\text{data}}(\mathcal{D}_{\theta}(s\cdot\mathcal{E}_{\theta}(\mathbf{x})),s\cdot\mathbf{x})+\mathcal{L}_{\text{lat.}}(\mathcal{E}_{\theta}(\mathbf{x}),\mathbf{z}_{\text{ref}}),$ (5)
@@ 3.3 Joint Neural Prior
Figure 2: The tSNE visualization (Maaten and Hinton, 2008) of the prior latent before and after prior convergence from augmented VCTK (Veaux and Yamagishi, 2017). Note the difference between the scale of axes of the two figures.
Table 1: Evaluation results on in-domain tasks, including GSR and denoising benchmarks. The test sets covers the test splits of used datasets, simulated unseen datasets, and in-the-wild data. Best is bolded, and second best is underlined. * denotes results taken from the original paper. VoiceBridge outperforms most baselines in nearly all metrics with a single inference step, demonstrating strong restoration ability in complex environments.
Task Dataset Data Domain Model PESQ($\uparrow$) SIG($\uparrow$) BAK($\uparrow$) OVRL($\uparrow$) UTMOS($\uparrow$) WVMOS($\uparrow$) NISQA($\uparrow$) General Speech Restoration Voicefixer-GSR In-Domain Data Input 1.943 2.962 2.858 2.395 2.633 1.965 2.674 Miipher 1.491 3.335 3.992 3.051 3.617 4.171 4.136 VoiceFixer 2.043 3.302 3.971 3.005 3.439 3.800 4.159 Resemble-Enhance 2.000 3.412 4.045 3.137 3.608 4.224 4.507 UniverSE++ 2.346 3.275 3.961 2.976 3.376 3.648 4.009 AnyEnhance - 3.406* 4.073* 3.136* - - 4.308* VoiceBridge 2.471 3.494 4.113 3.285 4.339 4.295 4.203 DNS-with- Reverb Out-Domain Data Input 1.176 1.767 1.481 1.377 1.323 -0.019 1.689 Miipher 1.324 3.401 3.975 3.128 2.973 3.195 3.629 VoiceFixer 1.407 3.355 3.982 3.052 2.526 2.961 3.704 Resemble-Enhance 1.154 3.502 4.021 3.205 2.756 3.248 4.583 UniverSE++ 1.098 2.548 3.649 2.244 1.669 2.463 3.006 AnyEnhance - 3.500* 4.040* 3.204* - - 3.722* VoiceBridge 1.583 3.581 4.127 3.318 3.382 4.195 4.588 DNS-Real Out-Domain Data (Real Data) Input - 2.985 2.510 2.212 1.940 1.483 2.160 Miipher - 3.325 3.976 3.171 3.911 3.171 4.124 VoiceFixer - 3.174 3.919 2.875 2.351 2.938 3.529 Resemble-Enhance - 3.395 3.993 3.100 2.767 3.298 4.329 UniverSE++ - 2.999 3.660 2.641 2.306 2.603 3.317 AnyEnhance - 3.488* 3.977* 3.161* - - - VoiceBridge - 3.473 4.025 3.185 4.223 4.045 4.501 Speech Enhancement VB-Demand In-Domain Data Input 1.972 3.343 3.124 2.694 3.063 2.985 2.998 SBSE 2.140 2.571 3.658 2.946 3.407 3.819 3.976 VoiceFixer 2.454 3.454 4.047 3.174 3.645 4.129 4.433 Resemble-Enhance 2.352 3.454 3.986 3.141 3.678 4.270 4.503 UniverSE++ 3.020 3.486 4.042 3.198 3.968 4.385 4.503 VoiceBridge 2.831 3.483 4.062 3.216 4.296 4.392 4.536 WSJ0-CHiME3 Out-Domain Data Input 1.248 2.486 1.868 1.795 1.823 0.128 1.520 SGMSE+ 1.723 3.517 3.946 3.191 3.282 3.547 4.294 StoRM 1.826 3.452 4.115 3.214 3.468 3.618 4.611 VoiceFixer 1.493 3.293 4.059 3.052 2.863 3.255 3.921 Resemble-Enhance 1.171 3.524 4.113 3.263 2.749 3.474 4.610 UniverSE++ 1.320 3.173 3.997 2.914 2.984 3.248 3.918 VoiceBridge 1.742 3.556 4.118 3.297 4.445 4.374 4.649
@@
$\mathcal{L}_{\text{np-enc}}$ $=\mathcal{L}^{\text{ep}}_{\text{data}}(\mathcal{D}(s\cdot\mathcal{E}^{\text{np}}_{\theta}(\mathbf{x}_{1})),s\cdot\hat{\mathbf{x}}_{0})+\mathcal{L}_{\text{lat.}}(\mathcal{E}^{\text{np}}_{\theta}(\mathbf{x}_{1}),\mathbf{z}_{0}),$ (6)
where $\mathcal{L}_{\text{lat.}}$ has been the joint neural prior loss consisting of MSE loss and cosine similarity loss as introduced in Appendix D, and $\mathcal{L}_{\text{data}}^{\text{ep}}$ is the data-space loss in EP-VAE training modified for prior convergence. Prior works (Fang et al., 2021) applies a KL convergence loss to a VAE for speech denoising. In contrast, we add objectives including cosine similarity in latent space and multiple supervision signals in the reconstructed data space, in order to reduce the distance between the latents in all geometrical measures. We prove our method achieves better results in Appendix I.
@@
Namely, as shown in Equation 6, the encoder $\mathcal{E}^{\text{np}}$ is fine-tuned with changed objectives to converge different LQ priors to HQ target, rather than still learning the latent presentation for reconstruction as Equation 5. Moreover, the EP constraint is preserved. When reducing the distance between LQ priors $\mathbf{z}_{1}$ and the HQ target $\mathbf{z}_{0}$ to achieve $\mathbf{z}_{1}^{\text{np}}$, scale equivariance is simultaneously required, which inherits the EP property to measure the consistency at varying energy levels. As shown in Figure 2, this technique significantly closes the distance between different LQ priors and the HQ target in the latent space, and naturally reduces the burden of LBM modeling. A further quantitative analysis on the prior distance is shown in Appendix H
3.4 From Denoiser to Generator